OpenStudy (anonymous):

What is the best approximation of the solution to the system to the nearest integer values? Graph of a system of linear equations. Equation 1 is x plus y equals 6. Equation 2 is x minus 6y equals negative 9. ( , )

OpenStudy (anonymous):

( , )

pooja195 (pooja195):

x+y=6 (isolate x) x=6-y <---equation 1 x-6y=-9 <---equation 2 we know what x is so 6-y-6y=-9 solve for y

OpenStudy (anonymous):

I'm Confused

pooja195 (pooja195):

On what part?

OpenStudy (anonymous):

how do I get y

pooja195 (pooja195):

ok \[\huge~\rm−y+6+y=6\] combine like terms -1+1= ?

OpenStudy (anonymous):

um.

pooja195 (pooja195):

it cancels out so you get 0

pooja195 (pooja195):

wait... does the answer have to be coordinates?

OpenStudy (anonymous):

its like the last one i think

pooja195 (pooja195):

hmm?

OpenStudy (anonymous):

wait (4,2)

OpenStudy (anonymous):

or (2,4)

pooja195 (pooja195):

im getting infinte solutions...

pooja195 (pooja195):

Wait i typed the problem wrong lol im sorry ok give me a min

OpenStudy (anonymous):

ok lol

pooja195 (pooja195):

(4,2) is correct

OpenStudy (anonymous):

Thank you!!

pooja195 (pooja195):

yw ^-^